OpenStudy (anonymous):

Let f(x) = 8x3 - 28x + 61 and g(x) = 2x + 5. Find f of x over g of x. What I have: 4x^2 2x + 5 / 8x^3 – 28x + 61 -8x^3 + 20x^2 4x^2(2x + 5) = 8x^3 + 20x^2

OpenStudy (anonymous):

@DanJS could you help me again?

OpenStudy (anonymous):

put one on top of the other, that is all

OpenStudy (anonymous):

?

OpenStudy (anonymous):

f of x over g of x is \[\frac{f(x)}{g(x)}\] that is all

OpenStudy (anonymous):

it wants me to actually solve it... not just show the equation...

OpenStudy (anonymous):

it is not an equation, there is nothing to "solve"

OpenStudy (anonymous):

\[f(x)=x-3,g(x)=x^2+1\\ \frac{f(x)}{g(x)}=\frac{x-3}{x^2+1}\]

OpenStudy (anonymous):

*facepalm* f(x)/g(x)